OpenStudy (anonymous):

write an equation. the width of a rectangle is 9 inches less that the length. the perimeter is 86 inches.

OpenStudy (anonymous):

eqauation is L-x W-9-x eqaution- 9-x+x=86?

OpenStudy (anonymous):

or 2x-9=86

OpenStudy (anonymous):

Let x be the length. width = x – 9 Formula of a perimeter: 2l + 2w = P So, the formula is 2(x)+2(x-9) = 86

OpenStudy (anonymous):

Right? I'm not sure.

OpenStudy (anonymous):

Solving: 2x + 2(x-9) = 86 2x + 2x - 18 = 86 4x = 86 +18 4x = 104 x = 26

OpenStudy (anonymous):

Substitute: 2(26) + 2(26-9) = 86 52 + 34 = 86 86 = 86